OpenStudy (anonymous):

What is the solution to this system of linear equations? 2x–15 = y x = y + 24

OpenStudy (anonymous):

What methods can you use?

OpenStudy (anonymous):

matrices?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

2x-y=15 x-y=24(Subtract them) -------- x=-9 put value of x in first equation -9-x=24 x=-33

OpenStudy (anonymous):

preferably in eq2

OpenStudy (anonymous):

y=-33

OpenStudy (anonymous):

u will be getting fooled by @Romero

OpenStudy (anonymous):

just for the headsup